Fermi National Laboratory


Accelerator Update for December 12, 2005 - December 14, 2005

Monday December 12

The day shift began with the Tevatron (TeV) shot setup on hold due to a P1 transfer line power supply (HV703) problem, with the Antiproton Source (Pbar) and Switchyard (SY) off due to shot setup, with Operations monitoring the Recycler stash, with NuMI off due to a Horn chiller problem, and with MiniBooNE taking beam.

Operations put the TeV through a dry squeeze at 8:43 AM.

Operations prepared NuMI for an access at 9:21 AM.

Operations prepared the Collider Detector at Fermilab (CDF) and D-Zero (D0) for access at 9:37 AM.

Operations prepared the TeV sector F0 for access at 10:48 AM so that operators could inspect the HV703 magnet. What they found was a ground fault, a movable ladder had been pushed up against the bus. Fortunately, the copper bus wasn't damaged, just the ladder.

Operations turned off Booster and MiniBooNE at 11:15 AM in preparation for a Booster access.

Everybody came out of TeV access and Operations began turning on the TeV at 12:35 PM.

Booster technicians had completed replacing all the power amplifier air filters and Operations began turning on the Booster at 1:55 PM.

At 2:53 PM, Operations reported that experts were investigating a problem with a TeV quadrupole (B0Q3) power supply.

Pbar began stacking at 2:54 PM.

Operations established beam to the Meson MCenter experiment at 3:20 PM.

At 4:19 PM, Operations reported that a Booster RF station (BRF19) was tripping off every couple of minutes. An expert investigated.

Power supply experts reported at 4:31 PM that in the process of investigating B0Q3, the problem cleared itself. They decided to leave the supply as it was.

Operations put the TeV through a dry squeeze at 4:49 PM.

Operations put the TeV into shot setup at 5:24 PM.

At 7:10 PM, the TeV suffered a sector F3 quench. Operators investigated and found that beam had been injected on top of other beam; the bunch had not incremented properly.

Operations prepared the Recycler to stash antiprotons at 7:44 PM. The stash ended at 8:18 PM and Pbar resumed stacking about five minutes later. MCenter resumed taking beam about the same time.

Operations put the TeV into shot setup at 10:16 PM.

At 11:39 PM, Operations reported that a TeV RF station (TRF4) had tripped off for the first of two times in twenty minutes. It reset both times.

Tuesday December 13

The midnight shift began with TeV shot setup on hold due to a TRF4 problem, with NuMI off due to a horn chiller problem, and with MiniBooNE taking beam.

At 12:10 AM, Operations reported that TRF4 had tripped off after the ninth antiproton transfer. Operators held off any further transfers, investigated, and then contacted an expert.

The Recycler prepared to stash at 12:36 AM.

Operations established store 4543 at 12:40 AM with an initial luminosity of 37.96E30.

MCenter resumed taking beam at 1:22 AM.

The stash ended and Pbar resumed stacking at 1:24 AM.

At 2:37 AM, the TeV suffered a sector E1 quench. Operators investigated, but could find no cause for the quench. They called in an expert.

Cryo system techs at CHL reported at 4:08 AM that there was a problem with a TeV sector E2 wet engine. They began work on the engine.

The TeV expert reported at 5:07 AM that though there was a problem with a control card for T:VE45, the quench was probably due to instability.

Operations prepared D0 for access at 5:15 AM.

At 7:55 AM, a water expert reported that the MI 62 LCW pumps were off due to low level. They began a hunt for a possible NuMI leak.

Also at 7:55 AM, Operations reported that TRF2 tripped off. An operator notified a RF expert.

The Recycler began preparing to stash at 8:15 AM. The stash ended and Pbar resumed stacking at 9:23 AM. MCenter resumed taking beam.

Pbar experts began a study period at 9:55 AM.

D0 came out of access 12:50 PM.

Operations put the TeV through a wet squeeze at 2:03 PM.

Operations put the TeV into shot setup at 2:42 PM.

At 3:41 PM, Operations reported that there was a leak in the Booster 95 LCW system; everywhere had been checked except the tunnel. Operations turned off MiniBooNE in preparation for a Booster access.

At 4:35 PM, operators found the Booster LCW leak on a B3 magnet at the beginning of the MI-8 transfer line. Water techs began repairs.

At 4:53 PM, the TeV suffered a sector A1 quench during the end of the shot setup.

Water techs reported at 5:40 PM that they had completed the Booster LCW repair. Operations began turning on Booster at 6 PM.

Pbar and MCenter resumed taking beam at 6:23 PM.

Operations put the TeV through a dry squeeze at 9:01 PM.

Operations prepared the Recycler to stash at 9:09 PM. The stash ended at 10:10 PM. Pbar and MCenter resumed taking beam.

At 11:01 PM, Operations reported that the TeV sector F17 collimator wasn't working correctly. (Experts investigated and eventually determined that the collimator didn't have a flat surface. They moved the collimator out of the beam and disabled it.)

Wednesday December 14

The midnight shift began with experts investigating the TeV collimator problem, with Operations monitoring the stash, with Pbar stacking, with NuMI off due to Horn chiller problems, and with MiniBooNE and MCenter taking beam.

Operations loaded the TeV with a 12x0 study-store at 1:29 AM.

Operations put the TeV into shot setup at 2:30 AM.

Operations established store 4549 at 4:47 AM with an initial luminosity of 48.33E30. Pbar and MCenter resumed taking beam.

The Recycler began preparing to stash at 6:53 AM. The stash ended. Pbar and MCenter were both taking beam by 7:49 AM.

The Plans for Wednesday

The plans are to stack, stash, and store. The Run Coordinator will allow Pbar experts to continue with studies.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 12/14/2005   email Fermilab